OpenStudy (anonymous):

3x=-18+4y 16y=58+5x use elimination

OpenStudy (anonymous):

line up all the variables, one equation under the other. i think its easier setting it up 3x=-18+4y (A) with 5x=-58+16y (B) under it. multiply (A) by negative four and add down. So (A) looks like -12x=72-16y. add it to other equation. -7x=14 is what u get. then solve for x which equals 2. then plug into one of the equations to get y which will be 6